ACCURATE AB INITIO POTENTIAL ENERGY SURFACE AND PREDICTED SPECTROSCOPIC PROP-ERTIES OF BeH_{2}, BeD_{2} and BeHD

\underline{HUI LI} and ROBERT J. LE ROY, Guelph-Waterloo Centre for Graduate Work in Chemistry and Biochemistry, University of Waterloo, Waterloo, Ontario N2L 3G1, Canada.

A three-dimensional potential energy surface for the ground electronic state of BeH_{2} has been constructed from more than 6800 symmetry-unique ab initio points calculated at the icMRCI/aug-cc-pV5Z level, and corrected for core electron electron correlation computed at the MR-ACPF/cc-pCV5Z level. Calculated spectroscopic constants of BeH_{2} and BeD_{2} are in excellent agreement with recent experimental results, and spectral constants for BeHD are predicted using the same potential surface. The effect of different interpolation methods on predicted potential function values and on the calculated level energies and spectroscopic constants has been examined. Preliminary results of work on determining a global analytic function for this three-dimensional potential energy surface based on symmetric and antisymmetric stretching coordinates will be described.